# Rethinking Channel Dependence for Multivariate Time Series Forecasting: Learning from Leading Indicators

### 31 Jan 2024 v1

@@ Abstract
Recently, channel-independent methods have achieved state-of-the-art performance in multivariate time series (MTS) forecasting. Despite reducing overfitting risks, these methods miss potential opportunities in utilizing channel dependence for accurate predictions. We argue that there exist locally stationary lead-lag relationships between variates, i.e., some lagged variates may follow the leading indicators within a short time period. Exploiting such channel dependence is beneficial since leading indicators offer advance information that can be used to reduce the forecasting difficulty of the lagged variates. In this paper, we propose a new method named LIFT that first efficiently estimates leading indicators and their leading steps at each time step and then judiciously allows the lagged variates to utilize the advance information from leading indicators. LIFT plays as a plugin that can be seamlessly collaborated with arbitrary time series forecasting methods. Extensive experiments on six real-world datasets demonstrate that LIFT improves the state-of-the-art methods by 5.5% in average forecasting performance.
@@ 1 Introduction
Multivariate time series (MTS) forecasting, one of the most popular research topics, is a fundamental task in various domains such as weather, traffic, and finance. An MTS consists of multiple channels (a.k.a., variates [^1] We use the terms “variate” and “channel” interchangeably.), where each channel is a univariate time series. Many MTS forecasting researches argue each channel has dependence on other channels. Accordingly, numerous approaches adopt channel-dependent (CD) strategies and jointly model multiple variates by advanced neural architectures, such as GNNs (Wu et al., 2020; Cao et al., 2020), MLPs (Chen et al., 2023; Ekambaram et al., 2023), CNNs (Wu et al., 2023), and Transformers (Zhou et al., 2021; Wu et al., 2021).
Unexpectedly, CD methods have been defeated by recently proposed channel-independent (CI) methods (Nie et al., 2023; Zhou et al., 2023; Li et al., 2023a; Xu et al., 2023) and even a simple linear model (Zeng et al., 2023). These CI methods seperately forecast each univariate time series based on its own historical values, instead of referring to other variates. While only modeling cross-time dependence, CI Transformers (Nie et al., 2023; Zhou et al., 2023) surprisingly outperform CD Transformers that jointly model cross-time and cross-variate dependence (Grigsby et al., 2021; Zhang & Yan, 2023). One reason is that existing CD methods lack prior knowledge about channel dependence and may encounter the overfitting issue (Han et al., 2023). This gives rise to an interesting question: is there any explicit channel dependence that is effective to MTS forecasting?
@@
  • Extensive experimental results on six real-world datasets demonstrate that LIFT significantly improves the state-of-the-art methods in both short-term and long-term MTS forecasting. Specifically, LIFT makes an average improvement of 7.6% over CI models and 4.3% over CD models. We also introduce a lightweight yet strong method LightMTS, which enjoys high parameter efficiency and achieves the best performance on popular Weather and Electricity datasets.
@@
Figure 2: Illustration of our key idea. In one case of test data, the leading indicators of $v_{3}$ are changed to $v_{2}$ and $v_{4}$, which lead by five and three steps, respectively. Accordingly, we shift $v_{2}$ and $v_{4}$ by the corresponding leading steps to keep them always aligned with $v_{3}$.
@@ 3.1 Overview
Training scheme
We train the parameters of the backbone and Lead-aware Refiner by the MSE loss between the predictions $\bm{\widetilde{\mathcal{X}}}_{t+1:t+H}$ and the ground-truth values $\bm{{\mathcal{X}}}_{t+1:t+H}$ of all the $C$ variates.
Technical challenges
Notably, it is non-trivial to leverage the lead-lag relationships due to issues of efficiency and noise. As Eq. (1) requires $\mathcal{O}(L)$, a brute-force estimation method that searches all possible $\delta$ in $\{1,\cdots,L\}$ requires $\mathcal{O}(L^{2})$ computations. Also, $\bm{S}^{(j)}_{t}$ can contain some irrelevant patterns from leaders which are noise to the lagged variate.
To tackle these issues, we implement the Lead Estimator by an efficient algorithm of $\mathcal{O}(L\log L)$ complexity. And we develop the Lead-aware Refiner by an Adaptive Frequency Mixer that adaptively generates frequency-domain filters for the leading indicators and mixes desirable frequency components according to the cross-correlations and variate states.
@@ 3.2 Lead Estimator
Given the normalized lookback window $\bm{{{X}}}_{t-L+1:t}$, the Lead Estimator first computes the cross-correlation coefficients between each pair of variate $i$ and variate $j$, based on an extension of Wiener–Khinchin theorem (Wiener, 1930) (see the detailed proof in Appendix A). Formally, we estimate the coefficients for all possible leading steps in $\{0,\cdots,L-1\}$ at once by the following equation:
@@ 3.3 Lead-aware Refiner
For a comprehensive understanding of leading indicators, it is noteworthy that the lead-lag patterns also depend on variate states. Different variates lie in their specific states with some intrinsic periodicities (or trends), e.g., solar illumination is affected by rains in the short term but keeps its daily periodicity irrelated to rains. The state of a variate may also change over time, exhibiting different correlation strengths with other variates, e.g., correlations between the traffic speeds of two adjacent roads are strong within peak hours but much weaker within off-peak hours. This inspires us to estimate the state of each variate and filter out noisy frequency components of the leading indicators.
@@
Intuitively, the preliminary predictions deserve more refinement from the leading indicators when the estimated correlation $\bm{{R}}_{t}^{(j)}$ is large. To filter out noise in $V^{(j)}$ and $\bm{U}^{(j)}$, we employ a filter factory to generate $1+2K$ frequency-domain filters as defined below:
@@
Then, we gather all refinement suggestions from $K$ leading indicators and mix all frequency components by:
@@ 3.4 Discussion
Reasoning why CD models show inferior performance. Many variates are unaligned with each other, while traditional models (e.g., Informer (Zhou et al., 2021)) simply mix multivariate information at the same time step. Thus they introduce outdated information from lagged variates which are noise and disturb predicting leaders. Though other models (e.g., Vector Auto-Regression (Giannone et al., 2010)) memorize CD from different time steps by static weights, they can suffer from overfitting since the leading indicators and leading steps vary over time.
LIFT can cooperate with arbitrary time series forecasting backbones. When combining LIFT with a CI backbone, we decompose MTS forecasting into two stages which focus on modeling cross-time correlations and cross-variate correlations, respectively. This scheme avoids introducing noisy channel dependence during the first stage and may reduce optimization difficulty compared with traditional CD methods. When combining LIFT with a CD backbone, we expect LIFT to refine the rough predictions with the actual observations of leading indicators in $\bm{S}_{t}^{(j)}$.
LIFT alleviates distribution shifts by dynamically selecting and shifting indicators. Existing normalization-based methods (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023b) handle distribution shifts of the statistical properties (e.g., mean and variance) in the lookback window and the horizon window. Our work is orthogonal to them as we take a novel investigation into a different kind of distribution shifts in channel dependence.
@@ 4 Lightweight MTS Forecasting with LIFT
Thanks to the flexibility of LIFT, we introduce a lightweight MTS forecasting method named LightMTS using a simple linear backbone (Zeng et al., 2023; Li et al., 2023a). Specifically, we first conduct instance normalization, and then the linear backbone takes the normalized lookback window to give rough predictions.
LightMTS involves $LH$ parameters for the backbone, $(C+L)N$ parameters for state estimation, and $(2NK^{2}+3NK+N+1)(\lfloor H/2\rfloor+1)+3(\lfloor H/2\rfloor+1)^{2}$ parameters for the Adaptive Frequency Mixer, where $N$ and $K$ are relatively small in practice (e.g., $N=8$ and $K=4$). As we do not model time series representations in the high-dimensional latent space, LightMTS is more lightweight than deep neural networks such as Transformers (Zhang & Yan, 2023) and CNNs (Wu et al., 2023). Empirical evidence is provided in Appendix D.1, where the parameter efficiency of LightMTS keeps similar to DLinear (Zeng et al., 2023).
@@ 5.1 Experimental Settings
Datasets
We conduct extensive experiments on six widely-used MTS datasets, including Weather (Zeng et al., 2023), Electricity (Wu et al., 2020), Traffic (Lai et al., 2018), Solar (Liu et al., 2023a), Wind (Liu et al., 2022), and PeMSD8 (Song et al., 2020). We provide the dataset details in Appendix C.1 and conduct experiments on other datasets, including Illness (Zeng et al., 2023) and ETT datasets (Zhou et al., 2021), in Appendix D.2.
Comparison Methods
As LIFT can cooperate with arbitrary time series forecasting backbones, we verify the effectiveness of LIFT with (i) two state-of-the-art CI models: PatchTST (Nie et al., 2023) and DLinear (Zeng et al., 2023); (ii) the state-of-the-art CD model: Crossformer (Zhang & Yan, 2023); (iii): a classic GNN-based CD model: MTGNN (Wu et al., 2020). We use these models to instantiate the backbone of LIFT, while we keep the same model hyperparameters for fair comparison. We tune other hyperparameters as far as possible, using their official implementations. Furthermore, we also include the baselines of PatchTST, such as FEDformer (Zhou et al., 2022) and Autoformer (Wu et al., 2021).
Setups
All of the methods follow the same experimental setup with the forecast horizon $H\in\{24,48,96,192,336,720\}$ for both short-term and long-term forecasting. We collect some baseline results reported by PatchTST to compare performance with LightMTS, where PatchTST has tuned the lookback length $L$ of FEDformer and Autoformer. For other methods, we set $L$ to $336$. We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics.
@@ 5.2 Performance Evaluation
Weather 24 0.091 0.121 0.089 0.117 0.104 0.152 0.090 0.126 0.090 0.136 0.087 0.126 0.090 0.128 0.089 0.126 6.1%
48 0.115 0.156 0.113 0.154 0.137 0.194 0.115 0.163 0.116 0.172 0.111 0.166 0.117 0.170 0.117 0.169 5.3%
96 0.152 0.199 0.148 0.199 0.176 0.237 0.146 0.204 0.146 0.212 0.147 0.211 0.157 0.216 0.154 0.222 4.1%
192 0.197 0.243 0.190 0.240 0.220 0.282 0.190 0.248 0.202 0.271 0.196 0.264 0.205 0.269 0.204 0.269 4.6%
336 0.249 0.283 0.238 0.281 0.265 0.319 0.243 0.292 0.245 0.306 0.250 0.302 0.258 0.312 0.260 0.314 2.5%
720 0.320 0.335 0.313 0.334 0.323 0.362 0.316 0.350 0.344 0.379 0.324 0.361 0.335 0.369 0.337 0.366 2.4%
Electricity 24 0.099 0.196 0.092 0.186 0.110 0.209 0.099 0.197 0.095 0.195 0.097 0.195 0.097 0.195 0.096 0.194 3.3%
48 0.115 0.210 0.113 0.209 0.125 0.223 0.113 0.209 0.116 0.216 0.114 0.214 0.116 0.215 0.114 0.213 2.9%
96 0.130 0.222 0.127 0.220 0.140 0.237 0.131 0.226 0.142 0.243 0.137 0.236 0.138 0.238 0.132 0.231 3.5%
192 0.148 0.240 0.147 0.239 0.153 0.249 0.148 0.242 0.159 0.259 0.158 0.258 0.160 0.261 0.152 0.251 2.2%
336 0.167 0.261 0.165 0.260 0.169 0.267 0.163 0.261 0.188 0.290 0.182 0.285 0.193 0.284 0.175 0.273 3.2%
720 0.202 0.291 0.196 0.291 0.203 0.301 0.198 0.295 0.258 0.349 0.216 0.314 0.242 0.327 0.228 0.315 5.4%
Traffic 24 0.317 0.233 0.300 0.214 0.371 0.267 0.347 0.255 0.483 0.273 0.392 0.246 0.402 0.260 0.392 0.259 7.1%
48 0.348 0.248 0.329 0.236 0.393 0.276 0.367 0.260 0.513 0.290 0.428 0.289 0.450 0.274 0.436 0.281 5.0%
@@
720 0.434 0.287 0.429 0.286 0.466 0.315 0.452 0.305 0.584 0.369 0.543 0.322 0.616 0.352 0.532 0.340 5.5%
Solar 24 0.094 0.159 0.086 0.146 0.133 0.219 0.093 0.150 0.082 0.134 0.080 0.137 0.076 0.140 0.068 0.125 12.4%
48 0.158 0.226 0.137 0.200 0.190 0.267 0.146 0.199 0.146 0.203 0.140 0.180 0.131 0.185 0.124 0.172 12.7%
96 0.178 0.232 0.175 0.229 0.222 0.291 0.185 0.240 0.179 0.245 0.172 0.236 0.171 0.238 0.163 0.226 6.9%
192 0.191 0.239 0.189 0.249 0.249 0.309 0.194 0.253 0.201 0.259 0.185 0.249 0.184 0.249 0.178 0.244 6.9%
336 0.199 0.251 0.196 0.259 0.269 0.324 0.198 0.260 0.213 0.258 0.213 0.250 0.191 0.253 0.190 0.247 6.3%
720 0.220 0.270 0.201 0.264 0.271 0.327 0.207 0.260 0.229 0.282 0.210 0.271 0.205 0.268 0.195 0.250 9.9%
Wind 24 0.137 0.179 0.131 0.175 0.151 0.198 0.136 0.182 0.122 0.173 0.122 0.173 0.124 0.172 0.124 0.170 3.4%
48 0.163 0.200 0.155 0.196 0.175 0.214 0.158 0.200 0.147 0.194 0.147 0.194 0.149 0.192 0.148 0.191 3.1%
96 0.186 0.216 0.175 0.213 0.197 0.230 0.177 0.214 0.172 0.217 0.171 0.214 0.170 0.211 0.169 0.208 3.6%
192 0.204 0.229 0.191 0.224 0.218 0.245 0.193 0.227 0.192 0.237 0.190 0.229 0.186 0.223 0.184 0.220 4.3%
336 0.216 0.239 0.202 0.234 0.233 0.258 0.204 0.237 0.209 0.252 0.204 0.242 0.195 0.233 0.192 0.227 4.9%
720 0.231 0.253 0.215 0.247 0.254 0.278 0.224 0.256 0.233 0.275 0.213 0.247 0.200 0.236 0.200 0.232 6.3%
PeMSD8 24 0.288 0.252 0.288 0.251 0.361 0.318 0.306 0.265 0.303 0.253 0.290 0.252 0.314 0.257 0.306 0.256 5.1%
48 0.367 0.285 0.363 0.285 0.475 0.378 0.386 0.303 0.346 0.275 0.342 0.274 0.357 0.281 0.356 0.279 5.3%
96 0.447 0.324 0.413 0.310 0.562 0.421 0.449 0.336 0.379 0.295 0.367 0.290 0.393 0.304 0.386 0.297 7.6%
192 0.517 0.346 0.463 0.334 0.611 0.443 0.502 0.364 0.415 0.315 0.404 0.304 0.440 0.333 0.429 0.324 7.6%
336 0.550 0.358 0.508 0.357 0.648 0.462 0.532 0.380 0.460 0.339 0.426 0.319 0.468 0.350 0.441 0.333 8.4%
720 0.637 0.407 0.562 0.380 0.748 0.519 0.601 0.422 0.518 0.380 0.462 0.335 0.511 0.379 0.484 0.342 11.8%
@@
LIFT makes an average improvement of 7.6% over PatchTST and DLinear on the six datasets. Notably, PatchTST and DLinear surpass Crossformer and MTGNN by a large margin on Weather, Electricity, and Traffic datasets, indicating the challenge of modeling channel dependence. Intriguingly, LIFT significantly improves CI backbones by an average margin of 4.7% on these challenging datasets, achieving the best performance in most cases. This confirms that the locally stationary lead-lag relationship as explicit channel dependence is effective in MTS forecasting.
@@
LIFT makes an average improvement of 3.4% over Crossformer and MTGNN on the six datasets. As CD backbones outperform CI ones on Solar, Wind, and PeMSD8, we conjecture that these three datasets have fewer distribution shifts in channel dependence, leading to fewer overfitting risks. Even though the CD backbones have benefited from channel dependence via complex neural architectures, LIFT can still refine their predictions, e.g., improving Crossformer by 4.5% on PeMSD8. This indicates that existing CD approaches, though employing rich parameters, cannot fully exploit the lead-lag relationships without prior knowledge about the dynamic variation of leading indicators and leading steps.
It is noteworthy that Crossformer utilizes the attention mechanism to model channel dependence and mainly mixes up information from the variates that show similarity at the same time step. However, as illustrated in Figure 1, the target variate may not be similar to its leading indicators at the same time step, to which Crossformer pays insufficient attention. Besides, MTGNN learns a static graph structure among variates on the training data and aggregates information within a fixed subset of variates. Thus MTGNN may well fail at those test data under different distributions of leading indicators or leading steps. In contrast, LIFT dynamically identifies and shifts the leading indicators, mitigating the distribution shifts before CD modeling. We visualize the distribution shifts of lead-lag relationships on the Weather dataset in Appendix D.3.
@@
Moreover, we compare the performance of LightMTS and all baselines on Weather, Electricity, and Traffic datasets. We use the baseline results from PatchTST with the prediction length in {96, 192, 336, 720}. As shown in Figure 4(a), LightMTS with a simple linear layer as its backbone still shows considerable performance among the state-of-the-art models. In particular, LightMTS surpasses PatchTST, the complex Transformer model, by 3.2% on Weather and 0.7% on Electricity. However, PatchTST significantly outperforms LightMTS on the Traffic dataset. As Traffic contains the greatest number of variates with complex temporal patterns, it requires a strong backbone to model the intricate cross-time dependence. Nevertheless, LightMTS is still the most competitive baseline on Traffic.

### 07 Apr 2024 v4

@@ Abstract
Recently, channel-independent methods have achieved state-of-the-art performance in multivariate time series (MTS) forecasting. Despite reducing overfitting risks, these methods miss potential opportunities in utilizing channel dependence for accurate predictions. We argue that there exist locally stationary lead-lag relationships between variates, i.e., some lagged variates may follow the leading indicators within a short time period. Exploiting such channel dependence is beneficial since leading indicators offer advance information that can be used to reduce the forecasting difficulty of the lagged variates. In this paper, we propose a new method named LIFT that first efficiently estimates leading indicators and their leading steps at each time step and then judiciously allows the lagged variates to utilize the advance information from leading indicators. LIFT plays as a plugin that can be seamlessly collaborated with arbitrary time series forecasting methods. Extensive experiments on six real-world datasets demonstrate that LIFT improves the state-of-the-art methods by 5.5% in average forecasting performance. Our code is available at https://github.com/SJTU-Quant/LIFT.
@@ 1 Introduction
Multivariate time series (MTS) forecasting, one of the most popular research topics, is a fundamental task in various domains such as weather, traffic, and finance. An MTS consists of multiple channels (a.k.a., variates [^1] We use the terms “variate” and “channel” interchangeably.), where each channel is a univariate time series. Many MTS forecasting researches argue each channel has dependence on other channels. Accordingly, numerous approaches adopt channel-dependent (CD) strategies and jointly model multiple variates by advanced neural architectures, including GNNs (Wu et al., 2020; Cao et al., 2020; Huang et al., 2023; Yi et al., 2023a), MLPs (Chen et al., 2023; Ekambaram et al., 2023; Wang et al., 2024a; Yi et al., 2023b), CNNs (Wu et al., 2023), Transformers (Zhou et al., 2021; Ni et al., 2023; Wang et al., 2024b; Liu et al., 2023a), and others (Wu et al., 2024; Shen et al., 2024; Jia et al., 2023; Fan et al., 2024).
Unexpectedly, CD methods have been defeated by recently proposed channel-independent (CI) methods (Nie et al., 2023; Lee et al., 2024; Zhou et al., 2023; Jin et al., 2024; Cao et al., 2024; Chen et al., 2024; Dai et al., 2024) and even a simple linear model (Zeng et al., 2023; Li et al., 2023a; Xu et al., 2023). These CI methods seperately forecast each univariate time series based on its own historical values, instead of referring to other variates. While only modeling cross-time dependence, CI Transformers (Nie et al., 2023; Zhou et al., 2023) surprisingly outperform CD Transformers that jointly model cross-time and cross-variate dependence (Grigsby et al., 2021; Zhang & Yan, 2023). One reason is that existing CD methods lack prior knowledge about channel dependence and may encounter the overfitting issue (Han et al., 2023). This gives rise to an interesting question: is there any explicit channel dependence that is effective to MTS forecasting?
@@
  • Extensive experimental results on six real-world datasets demonstrate that LIFT significantly improves the state-of-the-art methods in both short-term and long-term MTS forecasting. Specifically, LIFT makes an average improvement of 7.9% over CI models and 3.2% over CD models. We also introduce a lightweight yet strong method LightMTS, which enjoys high parameter efficiency and achieves the best performance on popular Weather and Electricity datasets.
@@
Figure 2: Illustration of our key idea. In one case of test data, $v_{1}$ no longer leads $v_{3}$. Instead, the leading indicators of $v_{3}$ are $v_{2}$ and $v_{4}$, which lead by five and three steps, respectively. An intuitive idea is to shift $v_{2}$ and $v_{4}$ by the corresponding leading steps to keep them always aligned with $v_{3}$.
@@ 3.1 Overview
Training scheme. We can jointly train the backbone and Lead-aware Refiner by the MSE between $\bm{\widetilde{\mathcal{X}}}_{t+1:t+H}$ and the ground truth $\bm{{\mathcal{X}}}_{t+1:t+H}$. Alternatively, given a pretrained and frozen backbone, we can precompute the preliminary predictions only once on training data, reducing the time of hyperparameter tuning and GPU memory occupation during training.
Technical challenges. Notably, it is non-trivial to leverage the lead-lag relationships due to issues of efficiency and noise. As Eq. (1) requires $\mathcal{O}(L)$, a brute-force estimation method that searches all possible $\delta$ in $\{1,\cdots,L\}$ requires $\mathcal{O}(L^{2})$ computations. Also, $\bm{S}^{(j)}_{t}$ can contain some irrelevant patterns from leaders which are noise to the lagged variate.
To tackle these issues, we implement the Lead Estimator by an efficient algorithm of $\mathcal{O}(L\log L)$ complexity. And we develop the Lead-aware Refiner by an Adaptive Frequency Mixer that adaptively generates frequency-domain filters and mixes desirable frequency components according to the cross-correlations and variate states.
@@ 3.2 Lead Estimator
Given the normalized lookback window $\bm{{{X}}}_{t-L+1:t}$, the Lead Estimator first computes the cross-correlation coefficients between each pair of variate $i$ and variate $j$, based on an extension of Wiener–Khinchin theorem (Wiener, 1930) (see details in Appendix A). Formally, we estimate the coefficients for all possible leading steps in $\{0,\cdots,L-1\}$ at once by the following equation:
@@ 3.3 Lead-aware Refiner
For a comprehensive understanding of leading indicators, it is noteworthy that the lead-lag patterns also depend on variate states. Different variates lie in their specific states with some intrinsic periodicities (or trends), e.g., solar illumination is affected by rains in the short term but keeps its daily periodicity. The state of a variate may also change over time, exhibiting different correlation strengths with other variates, e.g., correlations between the traffic speeds of two adjacent roads are strong within peak hours but much weaker within off-peak hours. Therefore, the variate states are informative signals that can guide us to filter out uncorrelated patterns.
@@
Intuitively, the preliminary predictions deserve more refinement from the leading indicators when the estimated correlation $\bm{{R}}_{t}^{(j)}$ is large. To filter signals in $V^{(j)}$ and $\bm{U}^{(j)}$, we employ a filter factory to generate $12K+$ 1 frequency-domain filters as defined below:
@@
Then, we gather information from $K$ leading indicators and mix the frequency components by:
@@ 3.4 Discussion
Reasoning why CD models show inferior performance. Many variates are unaligned with each other, while traditional models (e.g., Informer (Zhou et al., 2021)) simply mix multivariate information at the same time step. Consequently, they introduce outdated information from lagged variates which are noise and disturb predicting leaders. Though other models (e.g., Vector Auto-Regression (Giannone et al., 2010)) memorize CD from different time steps by static weights, they can suffer from overfitting issues since the leading indicators and leading steps vary over time.
LIFT can cooperate with arbitrary time series forecasting backbones. When combining LIFT with a CI backbone, we decompose MTS forecasting into two stages which focus on modeling time dependence and channel dependence, respectively. This scheme avoids introducing noisy channel dependence during the first stage and may reduce optimization difficulty compared with traditional CD methods. When combining LIFT with a CD backbone, we expect LIFT to refine the rough predictions with the actual observations of leading indicators in $\bm{S}_{t}^{(j)}$.
LIFT alleviates distribution shifts by dynamically selecting and shifting indicators. Existing normalization-based methods (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023b) handle distribution shifts of the statistical properties (e.g., mean and variance) in the lookback window and the horizon window. Our work is orthogonal to them as we take a novel investigation into a different kind of distribution shifts in channel dependence (see visualization in Appendix D.2).
@@ 4 Lightweight MTS Forecasting with LIFT
Thanks to the flexibility of LIFT, we introduce a lightweight MTS forecasting method named LightMTS, where a simple linear layer serves as a CI backbone. Following Li et al. (2023a), we conduct instance normalization before preliminary forecasting to alleviate distribution shifts.
As we do not learn representations in the high-dimensional latent space, LightMTS is more lightweight than popular CD models, including Transformers (Zhang & Yan, 2023; Liu et al., 2023a) and CNNs (Wu et al., 2023). Empirical evidence is provided in Appendix D.1, where the parameter efficiency of LightMTS keeps similar to DLinear Zeng et al. (2023).
@@ 5.1 Experimental Settings
Datasets. We conduct extensive experiments on six widely-used MTS datasets, including Weather (Zeng et al., 2023), Electricity (Wu et al., 2020), Traffic (Lai et al., 2018), Solar (Liu et al., 2023a), Wind (Liu et al., 2022), and PeMSD8 (Song et al., 2020). We provide the dataset details in Appendix C.1 and conduct experiments on more datasets in Appendix D.3.
Comparison Methods. As LIFT can incorporate arbitrary time series forecasting backbones, we verify the effectiveness of LIFT with (i) two state-of-the-art CI models: PatchTST (Nie et al., 2023) and DLinear (Zeng et al., 2023); (ii) the state-of-the-art CD model: Crossformer (Zhang & Yan, 2023); (iii) a classic CD model: MTGNN (Wu et al., 2020). We use them to instantiate the backbone of LIFT, while we keep the same model hyperparameters for fair comparison. We also include the baselines of PatchTST, such as FEDformer (Zhou et al., 2022) and Autoformer (Wu et al., 2021).
Setups. All of the methods follow the same experimental setup with the forecast horizon $H\in\{24,48,96,192,336,720\}$ for both short-term and long-term forecasting. We collect some baseline results reported by PatchTST to compare performance with LightMTS, where PatchTST has tuned the lookback length $L$ of FEDformer and Autoformer. For other methods, we set $L$ to $336$. We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics.
@@ 5.2 Performance Evaluation
Weather 24 0.091 0.122 0.089 0.119 0.104 0.152 0.090 0.125 0.086 0.126 0.086 0.126 0.090 0.128 0.090 0.126 4.7%
48 0.119 0.164 0.115 0.158 0.137 0.194 0.114 0.163 0.112 0.166 0.112 0.165 0.117 0.170 0.115 0.167 5.3%
96 0.152 0.199 0.146 0.196 0.176 0.237 0.145 0.203 0.145 0.209 0.146 0.210 0.157 0.216 0.154 0.212 5.0%
192 0.197 0.243 0.190 0.238 0.220 0.282 0.189 0.249 0.197 0.264 0.196 0.262 0.205 0.269 0.203 0.266 4.3%
336 0.249 0.283 0.243 0.281 0.265 0.319 0.243 0.292 0.246 0.309 0.245 0.305 0.258 0.312 0.256 0.308 3.0%
720 0.320 0.335 0.315 0.333 0.323 0.362 0.317 0.349 0.323 0.364 0.321 0.360 0.335 0.369 0.333 0.365 1.4%
Electricity 24 0.099 0.196 0.094 0.190 0.110 0.209 0.099 0.197 0.095 0.195 0.093 0.193 0.097 0.195 0.094 0.193 3.6%
48 0.115 0.210 0.110 0.205 0.125 0.223 0.113 0.209 0.116 0.216 0.113 0.211 0.116 0.215 0.112 0.211 4.0%
96 0.130 0.222 0.128 0.222 0.140 0.237 0.130 0.225 0.142 0.243 0.138 0.238 0.138 0.238 0.133 0.233 2.9%
192 0.148 0.240 0.147 0.239 0.153 0.249 0.148 0.242 0.159 0.259 0.154 0.251 0.160 0.261 0.153 0.252 2.7%
336 0.167 0.261 0.163 0.257 0.169 0.267 0.163 0.261 0.192 0.293 0.176 0.276 0.193 0.284 0.187 0.275 3.8%
720 0.202 0.291 0.195 0.289 0.203 0.301 0.198 0.295 0.264 0.353 0.224 0.312 0.242 0.327 0.216 0.305 6.6%
Traffic 24 0.323 0.235 0.300 0.214 0.371 0.267 0.347 0.255 0.483 0.273 0.392 0.246 0.402 0.260 0.392 0.259 7.3%
48 0.342 0.240 0.329 0.236 0.393 0.276 0.367 0.260 0.513 0.290 0.428 0.289 0.450 0.274 0.436 0.281 4.4%
@@
720 0.434 0.287 0.429 0.286 0.466 0.315 0.454 0.307 0.584 0.369 0.543 0.322 0.616 0.352 0.532 0.340 5.4%
Solar 24 0.095 0.160 0.087 0.147 0.133 0.219 0.093 0.149 0.082 0.134 0.079 0.129 0.070 0.125 0.069 0.122 11.0%
48 0.153 0.227 0.143 0.200 0.190 0.267 0.145 0.197 0.146 0.203 0.140 0.178 0.131 0.180 0.130 0.177 11.0%
96 0.176 0.227 0.174 0.224 0.222 0.291 0.185 0.238 0.179 0.245 0.174 0.224 0.167 0.224 0.166 0.223 6.2%
192 0.205 0.260 0.190 0.245 0.249 0.309 0.194 0.253 0.204 0.254 0.197 0.250 0.180 0.243 0.179 0.239 7.6%
336 0.200 0.252 0.194 0.249 0.269 0.324 0.198 0.260 0.216 0.257 0.204 0.254 0.191 0.251 0.190 0.245 7.5%
720 0.229 0.282 0.203 0.261 0.271 0.327 0.207 0.260 0.211 0.250 0.203 0.256 0.197 0.256 0.195 0.251 8.5%
Wind 24 0.137 0.179 0.131 0.175 0.151 0.198 0.136 0.182 0.122 0.173 0.121 0.168 0.124 0.172 0.124 0.170 3.8%
48 0.163 0.200 0.155 0.196 0.175 0.214 0.159 0.200 0.147 0.194 0.147 0.189 0.149 0.192 0.148 0.191 3.4%
96 0.186 0.216 0.175 0.213 0.197 0.230 0.177 0.214 0.172 0.218 0.169 0.208 0.170 0.211 0.169 0.208 4.1%
192 0.204 0.229 0.191 0.224 0.218 0.245 0.193 0.226 0.189 0.230 0.187 0.222 0.186 0.223 0.184 0.220 4.4%
336 0.216 0.239 0.202 0.234 0.233 0.258 0.205 0.238 0.201 0.240 0.199 0.232 0.195 0.233 0.192 0.227 4.6%
720 0.231 0.253 0.215 0.247 0.254 0.278 0.225 0.256 0.237 0.286 0.224 0.254 0.200 0.236 0.200 0.232 6.0%
PeMSD8 24 0.289 0.247 0.285 0.246 0.361 0.318 0.306 0.265 0.303 0.253 0.290 0.252 0.314 0.257 0.306 0.256 5.2%
48 0.367 0.281 0.356 0.277 0.475 0.378 0.386 0.303 0.346 0.275 0.342 0.274 0.357 0.281 0.356 0.279 5.8%
96 0.445 0.316 0.410 0.305 0.562 0.421 0.449 0.336 0.379 0.295 0.367 0.290 0.393 0.304 0.386 0.297 7.6%
192 0.519 0.354 0.471 0.337 0.611 0.443 0.502 0.364 0.415 0.315 0.404 0.304 0.440 0.333 0.429 0.324 7.6%
336 0.562 0.366 0.511 0.353 0.648 0.462 0.532 0.379 0.460 0.339 0.426 0.319 0.468 0.350 0.441 0.333 9.1%
720 0.653 0.403 0.563 0.378 0.748 0.519 0.597 0.414 0.518 0.380 0.462 0.335 0.511 0.379 0.484 0.342 12.3%
@@
LIFT makes an average improvement of 7.9% over PatchTST and DLinear on the six datasets. Notably, PatchTST and DLinear surpass Crossformer and MTGNN by a large margin on Weather, Electricity, and Traffic datasets, indicating the challenge of modeling channel dependence. Intriguingly, LIFT significantly improves CI backbones by an average margin of 4.7% on these challenging datasets, achieving the best performance in most cases. This confirms that LIFT can reduce overfitting risks by introducing prior knowledge about channel dependence.
@@
LIFT makes an average improvement of 3.2% over Crossformer and MTGNN on the six datasets. As CD backbones outperform CI ones on Solar, Wind, and PeMSD8, we conjecture that these datasets have fewer distribution shifts in channel dependence, leading to fewer overfitting risks. Even though the CD backbones have benefited from channel dependence, LIFT can still refine their predictions, e.g., improving Crossformer by 4.5% on PeMSD8. This indicates that existing CD approaches cannot fully exploit the lead-lag relationships without prior knowledge about the dynamic variation of leading indicators and leading steps. Moreover, Crossformer mixes information from the variates that show similarity at the same time step but pays insufficient attention to the different yet informative signals of leading indicators. MTGNN learns a static graph structure among variates on the training data and aggregates information within a fixed subset of variates. MTGNN may well suffer from distribution shifts in channel dependence, while LIFT dynamically selects leading indicators and reduces overfitting risks.
@@
Moreover, we compare the performance of LightMTS and all baselines on Weather, Electricity, and Traffic datasets. We borrow the baseline results from the paper of PatchTST with $H\in\{96,192,336,720\}$. As shown in Figure 4(a), LightMTS with a simple linear layer as its backbone still shows considerable performance among the state-of-the-art models. In particular, LightMTS surpasses PatchTST, the complex Transformer model, by 3.2% on Weather and 0.7% on Electricity. However, PatchTST significantly outperforms LightMTS on the Traffic dataset. As Traffic contains the greatest number of variates with complex temporal patterns, it requires a strong backbone to model the intricate cross-time dependence. Nevertheless, LightMTS is still the most competitive baseline on Traffic.